# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v5

@@ III-B Feedback DRL control hybrid scheme
In contrast to the comprehensive approach of end-to-end policies discussed in Section II, which excels in handling versatile locomotion skills and complex terrains with minimal interface times, the Feedback DRL Control Hybrid Scheme integrates DRL policies as LL controllers. These LL controllers, replacing traditional model-based feedback mechanisms, work in conjunction with HL planners that process terrain information, plan future walking paths, and maintain robust locomotion stability.
@@
For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [58]. Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness. A more dynamic approach involves online planning, which has shown greater adaptability and efficiency. One notable framework combines a conventional foot planner with an LL DRL policy [38], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands. Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies. However, this work has not yet been transferred from simulation to real-world applications. Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [39]. This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
@@
A recent development has shown that focusing solely on foot placement might restrict the stability and adaptability of locomotion, particularly in complex maneuvers. A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [61]. This system employs a residual learning architecture, where the DRL policy’s outputs are merged with the planner’s directives before being relayed to the PD controller. This integrated approach not only concerns itself with foot placement but also generates comprehensive trajectories for trunk position, orientation, and ankle yaw angle, enabling the robot to perform a wide array of locomotion skills including walking, squatting, turning, and stair climbing.
@@
Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities. These policies have demonstrated high efficiency in locomotion and accurate reference tracking [62]. Despite their extensive capabilities, such policies generally require short inference times, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited. Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [62].
@@
Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [39]. This capability effectively addresses the navigational limitations observed in end-to-end frameworks. Moreover, unlike the deep planning hybrid scheme where modifications post-policy establishment can be cumbersome, this hybrid scheme offers enhanced flexibility for on-the-fly adjustments.
@@
Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models. Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation and hinder rapid deployment and iterative enhancements [62].
@@ III-C Learned hierarchy framework
The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [25]. In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance. An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths. This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal. The framework was later enhanced to include imitation learning, facilitating the replication of dynamic human-like movements within the simulation environment [26].
@@
However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks have predominantly been validated only in simulations. The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
@@
Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [41].
@@
To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [59]. It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as a LL tracking controller. Additionally, binocular RGB vision input enhances perception, enabling precise loco-manipulation and dynamic locomotion tasks such as jumping, walking, folding clothes, and rearranging objects. This shadowing-based imitation learning approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
@@
Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics. Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [25]. As discussed in the previous section, the benefits of integrating RL-based planners with RL-based controllers have been effectively demonstrated. This combination enables the framework to adeptly manage a diverse array of environments and tasks.
@@
Within such a framework, HL policy is optimised for strategic planning and achieving specific goals. This optimisation allows for targeted enhancements depending on the tasks at hand. Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives. Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
@@
Additionally, the training process for each policy within the hierarchy demands considerable computational resources [41]. The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions. This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
@@ IV Limitations and Challenges
The end-to-end and hierarchical frameworks detailed in Sections II and  III represent the state-of-the-art in DRL-based bipedal locomotion, demonstrating significant capabilities on specific tasks. However, a significant gap persists between these specialised successes and the realisation of a truly unified framework as defined in Section  I.
Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these criteria. The path to a unified framework is hindered by several interconnected issues. At a foundational level, a primary difficulty involves the limitations and challenges in achieving both generalisation and precision (Section IV-A). This is further complicated by the practical barrier of the ‘reality gap’ in transferring policies from simulation to physical robots (Section IV-B). Ultimately, these issues culminate in the critical challenges of ensuring safety and interpretability for robust deployment in real-world, safety-critical situations (Section IV-C).
@@ IV-A Generalisation and precision
A central challenge in applying DRL to bipedal locomotion is the need to simultaneously achieve high generalisation across diverse skills and transverse to all kinds of terrains, and high precision in specific tasks. This remains a fundamental obstacle to realising truly unified and capable frameworks.
This capability gap is evident in the current literature. Many approaches excel at generalisation, demonstrating policies that enable versatile skills like walking and jumping [44, 35] and can transfer to different terrains [49, 33, 63]. However, these generalized policies often lack the fidelity required for high-precision tasks, such as exact foot placement [38, 33, 54] or maintaining a specific velocity with minimal error [52]. Conversely, controllers specialized for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [48], yet they cannot generalize these capabilities to a broader range of tasks. Thus, the development of a single, unified framework that concurrently exhibits both broad competency and high fidelity remains a largely unresolved challenge.
@@
  • Limited terrain and gait patterns: The failure to generalise is often a direct result of training on insufficiently diverse environments or with a restricted set of behaviours. Models trained on limited terrain are brittle when faced with novel surfaces, while a limited gait pattern library prevents them from adapting to tasks requiring new motor skills.
@@
  • Poor command tracking: The learning signals for generalisation and precision are often in direct conflict. generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations. In contrast, precision demands restrictive signals that minimize command tracking error. Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [34, 35].
@@
  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [40, 64, 65]. This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only provides support for the success of the final task, thus the task, hence the community’s frequent resort to reward shaping and curriculum learning [53, 32, 58, 66]. Consequently, the immense amount of data required for an agent to explore, stumble upon a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation just to make training tractable [67, 68, 25, 26].
@@
  • High-quality data scarcity: As highlighted in our discussion in Section LABEL:subsec:guided_learning, the scarcity of high-quality demonstrations is a key bottleneck.Such data provides essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while bypassing extensive, unsafe exploration [26, 25, 69]. This scarcity stems from the fact that scalable human data is difficult to transfer due to embodiment mismatch [70, 71, 59], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [20, 21].
@@
These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly embodied in the designs of the field’s dominant control architectures. End-to-end frameworks attempt a holistic solution, learning a single, monolithic policy that must implicitly resolve all challenges simultaneously. While this approach can yield highly versatile and dynamic behaviours [63], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within the system’s unstable dynamics. This often results in a notable lack of the fidelity and precision that hierarchical systems can enforce [35]. Conversely, hierarchical frameworks, for example, are a direct architectural response to the lack of skill compositionality. By employing a “divide and conquer” strategy, they use a HL policy to sequence a library of LL, often model-based, controllers. This structure is designed to enforce precision and manage the complex dynamics at a lower level [21, 54, 39]. However, as noted, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [38].
IV-B Reality gap: From simulation to the physical world

### v6

@@ III-B Feedback DRL control hybrid scheme
In contrast to the comprehensive approach of end-to-end policies discussed in Section II, which excels in handling versatile locomotion skills and complex terrains with minimal inference time, the Feedback DRL Control Hybrid Scheme integrates DRL policies as LL controllers. These LL controllers, replacing traditional model-based feedback mechanisms, work in conjunction with HL planners that process terrain information, plan future walking paths, and maintain robust locomotion stability.
@@
For instance, gait libraries, which provide predefined movement references based on user commands, have been integrated into such frameworks [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real]. Despite the structured approach of using gait libraries, their static nature offers limited adaptability to changing terrains, diminishing their effectiveness. A more dynamic approach involves online planning, which has shown greater adaptability and efficiency. One notable framework combines a conventional foot planner with an LL DRL policy [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P], delivering targeted footsteps and directional guidance to the robot, thereby enabling responsive and varied walking commands. Moreover, HL controllers can provide additional feedback to LL policies, incorporating CoM or end-feet information, either from model-based methods or other conventional control strategies. However, this work has not yet been transferred from simulation to real-world applications. Later, a similar structure featuring an HL foot planner and an LL DRL-based policy was proposed [2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]. This strategy not only achieved a successful sim-to-real transfer but also enabled the robot to navigate omnidirectionally and avoid obstacles.
@@
A recent development has shown that focusing solely on foot placement might restrict the stability and adaptability of locomotion, particularly in complex maneuvers. A new framework integrates a model-based planner with a DRL feedback policy to enhance bipedal locomotion’s agility and versatility, displaying improved performance [2024_Li_model_based_LLDRL_hybrid_control_scheme]. This system employs a residual learning architecture, where the DRL policy’s outputs are merged with the planner’s directives before being relayed to the PD controller. This integrated approach not only concerns itself with foot placement but also generates comprehensive trajectories for trunk position, orientation, and ankle yaw angle, enabling the robot to perform a wide array of locomotion skills including walking, squatting, turning, and stair climbing.
@@
Compared to traditional model-based controllers, learned DRL policies provide a comprehensive closed-loop control strategy that does not rely on assumptions about terrain or robotic capabilities. These policies have demonstrated high efficiency in locomotion and accurate reference tracking [quadRLLL]. Despite their extensive capabilities, such policies generally require short inference time, making DRL a preferred approach in scenarios where robustness is paramount or computational resources on the robot are limited. Nonetheless, these learning algorithms often face challenges in environments characterised by sparse rewards, where suitable footholds like gaps or stepping stones are infrequent [quadRLLL].
@@
Additionally, an HL planner can process critical data such as terrain variations or obstacles and generate precise target locations for feet or desired walking paths, instead of detailed terrain data, which can significantly expedite the training process [2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]. This capability effectively addresses the navigational limitations observed in end-to-end frameworks. Moreover, unlike the deep planning hybrid scheme where modifications post-policy establishment can be cumbersome, this hybrid scheme offers enhanced flexibility for on-the-fly adjustments.
@@
Despite the significant potential demonstrated by previous studies, integrating DRL-based controllers with sophisticated and complex HL planners still presents limitations compared to more integrated frameworks such as end-to-end and deep planning models. Specifically, complex HL model-based planners often require substantial computational resources to resolve problems, rely heavily on model assumptions, necessitate extensive training periods, demand large datasets for optimisation, and hinder rapid deployment and iterative enhancements [quadRLLL].
@@ III-C Learned hierarchy framework
The genesis of this framework was within a physics engine, aimed at validating its efficiency through simulation [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.]. In this setup, LL policies, informed by human motions or trajectories generated via TO, strive to track these trajectories as dictated by the HL planner while maintaining balance. An HL policy is then introduced, pre-trained with long-term task goals, to navigate the environment and identify optimal paths. This structure enabled sophisticated interactions such as guiding a biped to dribble a soccer ball towards a goal. The framework was later enhanced to include imitation learning (IL), facilitating the replication of dynamic human-like movements within the simulation environment [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
@@
However, despite its structured and layered approach, which allows for the reuse of learned behaviours to achieve long-term objectives, these frameworks has been validated only in simulations. The interface designed manually between the HL planner and the LL controller sometimes leads to suboptimal behaviours, including stability issues like falling.
@@
Expanding the application of this framework, a sim-to-real strategy for a wheeled bipedal robot was proposed, focusing the LL policy on balance and position tracking, while the HL policy enhances safety by aiding in collision avoidance and making strategic decisions based on the orientation of subgoals [2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot].
@@
To fully leverage its potential, HumanPlus has been developed as a versatile framework for humanoid robots, integrating hierarchical learning, multimodal perception, and real-world imitation [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation]. It employs a two-layer structure, where HIT learns from human demonstrations, trained on AMASS, and HST acts as an LL tracking controller. Additionally, binocular RGB vision input enhances perception, enabling precise loco-manipulation and dynamic locomotion tasks such as jumping, walking, folding clothes, and rearranging objects. This shadowing-based IL approach improves adaptability, making it a promising framework for transferring human-like skills to robots.
@@
Learning complex locomotion skills, particularly when incorporating navigation elements, presents a significant challenge in robotics. Decomposing these tasks into distinct locomotion and navigation components allows robots to tackle more intricate activities, such as dribbling a soccer ball [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer.]. As discussed in the previous section, the benefits of integrating RL-based planners with RL-based controllers have been effectively demonstrated. This combination enables the framework to adeptly manage a diverse array of environments and tasks.
@@
Within such a framework, the HL policy is optimised for strategic planning and achieving specific goals. This optimisation allows for targeted enhancements depending on the tasks at hand. Moreover, the potential for continuous improvement and adaptation through further training ensures that the system can evolve over time, improving its efficiency and effectiveness in response to changing conditions or new objectives. Despite the theoretical advantages, the practical implementation of this type of sim-to-real application for bipedal robots remains largely unexplored.
@@
Additionally, the training process for each policy within the hierarchy demands considerable computational resources [2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot]. The intensive nature of this training can lead to a reliance on the simulation environment, potentially causing the system to overfit to specific scenarios and thereby fail to generalise to real-world conditions. This limitation highlights a significant hurdle that must be addressed to enhance the viability of learned hierarchy frameworks in practical applications.
@@ IV Limitations and Challenges
The end-to-end and hierarchical frameworks detailed in Sections II and III represent the state of the art in DRL-based bipedal locomotion, demonstrating remarkable capabilities on specific tasks. However, a substantial gap remains between these task-oriented successes and the broader goal of achieving generalisation and adaptability across diverse morphologies, tasks, and environments. Bridging this gap requires more than incremental improvements—it demands the establishment of a unified framework that consolidates interfaces, training conventions, and evaluation protocols to systematically address the underlying limitations of current DRL pipelines.
As outlined in the following sections, the core challenges underlying this gap can be grouped into three interrelated aspects. At a foundational level, a primary difficulty involves the limitations and challenges in achieving both generalisation and precision (Section IV-A). This is further complicated by the practical barrier of the sim-to-real gap in transferring policies from simulation to physical robots (Section IV-B). Ultimately, these issues culminate in the critical challenges of ensuring safety and interpretability for robust deployment in real-world, safety-critical situations (Section IV-C).
@@ IV-A Generalisation and precision
A central challenge in applying DRL to bipedal locomotion is the need to simultaneously achieve high generalisation across diverse skills and traverse to all kinds of terrains, and high precision in specific tasks. This remains a fundamental obstacle to realising truly unified and capable frameworks.
This capability gap is evident in the current literature. Many approaches excel at generalisation, demonstrating policies that enable versatile skills such as walking and jumping [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining] and can transfer to different terrains [2021_siekmann_blind_DRL_stair, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_radosavovic_realworld_transformer_NNmodel]. However, these generalised policies often lack the fidelity required for high-precision tasks such as exact foot placement [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation] or maintaining a specific velocity with minimal error [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction]. Conversely, controllers specialised for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [2023_UCB_Cassie_RL_Jumping], yet they cannot generalise these capabilities to a broader range of tasks. Thus, the development of a single unified framework that concurrently exhibits both broad competency and high fidelity remains largely unresolved.
@@
  • Limited terrain and gait patterns: The failure to generalise is often a direct result of training on insufficiently diverse environments or with a restricted set of behaviours. Models trained on limited terrain are brittle when faced with novel surfaces, while a limited gait pattern library prevents adaptation to tasks requiring new motor skills.
@@
  • Poor command tracking: The learning signals for generalisation and precision are often in direct conflict. Generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations, whereas precision demands restrictive signals that minimise command-tracking error. Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
@@
  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [2017_DRL_survey, 2017_Schulman_PPO_instructure, 2023_Omur_comparison_RLalgorithm_bipedallocomotion]. This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only supports the success of the final task. Consequently, the immense amount of data required for an agent to explore, discover a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation merely to make training tractable [2021_Sim_Isaac, 2017_heess_emergence_animation_PPO_comprarisonrlalgorithm_differenttarrains, 2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task].
@@
  • High-quality data scarcity: As highlighted in Section II-A2, the scarcity of high-quality demonstrations is a key bottleneck. Such data provide essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while avoiding unsafe exploration [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2020_motioncapture_referencebased_e2e_simulation_pybullet_augmentedrewarddesign]. This scarcity stems from the difficulty of transferring scalable human data due to embodiment mismatch [2017_Ayusawa_motionretarget_optimization, 2019_iit_wholebody_retargeting, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy].
@@
These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly reflected in the design of the field’s dominant control architectures. End-to-end frameworks attempt a holistic solution, learning a single monolithic policy that must implicitly resolve all challenges simultaneously. While this approach can yield highly versatile and dynamic behaviours [2023_radosavovic_realworld_transformer_NNmodel], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within unstable system dynamics. This often results in a lack of the fidelity and precision that hierarchical systems can enforce [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining]. Conversely, hierarchical frameworks are a direct architectural response to the lack of skill compositionality. By employing a “divide and conquer” strategy, they use an HL policy to sequence a library of LL, often model-based, controllers. This structure enforces precision and manages complex dynamics at a lower level [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]. However, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P].
IV-B Challenges in transferring from simulation to reality